Drawing a Card. If one card is drawn from a deck, find the probability of getting these results Getting a queen


1
Expert's answer
2021-12-09T10:29:14-0500

The number of the queens in a deck= 4

If we draw one card at random from the 52 cards,

Then the probability of getting a queen

"=\frac{ Number\>of\>queens}{Number\>of\>all\>the\>cards}"


"=\frac{4}{52}=\frac{1}{13}"